# A Bayesian approach to functional regression: theory and computation

### v2

@@ Abstract
We propose a novel Bayesian methodology for inference in functional linear and logistic regression models based on the theory of reproducing kernel Hilbert spaces (RKHS’s). We introduce general models that build upon the RKHS generated by the covariance function of the underlying stochastic process, and whose formulation includes as particular cases all finite-dimensional models based on linear combinations of marginals of the process, which can collectively be seen as a dense subspace made of simple approximations. By imposing a suitable prior distribution on this dense functional space we can perform data-driven inference via standard Bayes methodology, estimating the posterior distribution through reversible jump Markov chain Monte Carlo methods. In this context, our contribution is two-fold. First, we derive a theoretical result that guarantees posterior consistency, based on an application of a classic theorem of Doob to our RKHS setting. Second, we show that several prediction strategies stemming from our Bayesian procedure are competitive against other usual alternatives in both simulations and real data sets, including a Bayesian-motivated variable selection method.
@@ 1 Introduction
The problem of predicting a scalar response from a functional covariate is one that has gained traction over the last few decades, as more and more data is being generated with an ever-increasing level of granularity in the measurements. While in principle the functional data could simply be regarded as a discretized vector in a very high dimension, there are often many advantages in taking into account its functional nature, ranging from modeling the correlation among points that are close in the domain, to extracting information that may be hidden in the derivatives of the function in question. As a consequence, numerous proposals have arisen on how to suitably deal with functional data, all of them encompassed under the term Functional Data Analysis (FDA), which essentially explores statistical techniques to process, model and make inference on data varying over a continuum. A partial survey on such methods is Cuevas (2014) or Goia and Vieu (2016), while a more detailed exposition of the theory and applications can be found for example in Hsing and Eubank (2015).
@@
FDA is undoubtedly an active area of research, which finds applications in a wide variety of fields, such as biomedicine, finance, meteorology or chemistry (Ullah and Finch, 2013). Accordingly, there are many recent contributions on how to tackle functional data problems, both from a theoretical and practical standpoint. Chief among them is the approach of reducing the problem to a finite-dimensional one, for example using a truncated basis expansion or spline interpolation methods (Müller and Stadtmüller, 2005; Aguilera and Aguilera-Morillo, 2013). At the same time, much effort has been put into building a sound theoretical basis for FDA, generalizing different frequentist concepts to the infinite-dimensional framework. Examples of this endeavor include the definition of centrality measures and depth-based notions for functional data (López-Pintado and Romo, 2009), functional ANOVA tests (Cuevas et al., 2004), a functional Mahalanobis distance (Galeano et al., 2015; Berrendero et al., 2020), or an extension of Fisher’s discriminant analysis for random functions (Shin, 2008), among many others. As the name suggests, FDA techniques are heavily inspired by functional analysis tools and methods: Hilbert spaces, linear operators, orthonormal bases, and so on. Incidentally, a notion that also intersects with the classical theory of machine learning and pattern recognition, and that has attained popularity in recent years, is that of reproducing kernel Hilbert spaces (RKHS’s) and their applications in functional data problems (e.g. Kupresanin et al., 2010; Yuan and Cai, 2010; Berrendero et al., 2018). On the other hand, Bayesian inference methods are ubiquitous in the realm of statistics, and though they also make use of random functions, the approach is slightly different from the FDA case. While there are recent works that offer a Bayesian treatment of functional data (e.g. Crainiceanu and Goldsmith, 2010; Shi and Choi, 2011), there is still no systematic approach to Bayesian methodologies within FDA. It is precisely at this relatively unexplored intersection between FDA and Bayesian methods that our work is aimed.
@@
In particular, our goal is to study functional regression problems, the infinite-dimensional equivalents of the typical statistical regression problems, from a Bayesian perspective. We follow the path started by Ferguson (1974) of setting a prior distribution on a functional space and using the corresponding posterior for inference. In our case, we use a particular RKHS as the ambient space, resulting in functional regression models that allow for a simple yet efficient Bayesian treatment of functional data. Moreover, we also study the basic theoretical question of posterior consistency in these RKHS models within the proposed Bayesian framework. Consistency and posterior concentration are a type of frequentist validation criteria that have arguably been an active point of research in the last few decades, particularly in infinite-dimensional settings (Amewou-Atisso et al., 2003; Choi and Ramamoorthi, 2008), and also in the functional regression case (Lian et al., 2016; Abraham and Grollemund, 2020). To put it simply, posterior consistency ensures that with enough data points, the Bayesian updating mechanism works as intended and the posterior distribution eventually concentrates around the true value of the parameters, supposing the model is well specified. We leverage the properties of RKHS’s and existing techniques to show that posterior consistency holds in our models under some mild identifiability conditions, thus providing a coherent background to our Bayesian approach.
@@
Finally, this theoretical side is complemented by extensive experimentation that showcases the application of the models in various prediction tasks. Following recent trends in Bayesian computation techniques, the posterior distribution is approximated via Markov chain Monte Carlo (MCMC) methods, specifically the reversible jump variant (RJMCMC) proposed by Green (1995). This computational work highlights the predictive performance of the proposed functional regression models, especially when compared with other usual frequentist methods.
$\bm{L^{2}}$-models, shortcomings and alternatives
@@
[…]
where $\alpha_{0}\in\mathbb{R}$, $\varepsilon$ is a random error term independent from $X$ with $\mathbb{E}[\varepsilon]=0$, and the functional slope parameter $\beta=\beta(\cdot)$ is a member of the infinite-dimensional space $L^{2}[0,1]$. In this case, the inference on $\beta$ is hampered by the fact that $L^{2}[0,1]$ is an extremely broad space that contains many non-smooth or ill-behaved functions, so any estimation procedure involving optimization on it will typically be hard. In spite of this, model (1.1) is not flexible enough to include “simple” finite-dimensional models based on linear combinations of the marginals, such as $Y=\alpha_{0}+\beta_{1}X(t_{1})+\cdots+\beta_{p}X(t_{p})+\varepsilon$ for some constants $\beta_{j}\in\mathbb{R}$ and instants $t_{j}\in[0,1]$, which are especially appealing to practitioners confronted with functional data problems; see Berrendero et al. (2024) for additional details on this. Moreover, the non-invertibility of the covariance operator associated with $X$, which plays the role of the covariance matrix in the infinite case, invalidates the usual least squares theory (Cardot and Sarda, 2018). Thus, some regularization or dimensionality reduction technique is needed for parameter estimation; see Reiss et al. (2017) for a summary of several widespread methods.
@@
$\mathbb{P}(Y=1\mid X)=\frac{1}{1+\exp\{-\alpha_{0}-\left\langle X,\beta\right\rangle\}},$ (1.2)
where $\alpha_{0}\in\mathbb{R}$ and $\beta\in L^{2}[0,1]$. In this situation, the most common way of estimating the slope function $\beta$ is via its maximum likelihood estimator (MLE). The same complications as in the linear case apply here, with the additional problem that in functional settings the MLE does not exist with probability one under fairly general conditions (see Berrendero et al., 2023).
It turns out that in both scenarios a natural alternative to the $L^{2}$-model is the so-called reproducing kernel Hilbert space (RKHS) model, which instead assumes the unknown functional parameter $\beta$ to be a member of the RKHS associated with the covariance function of the process $X$. As we will show later on, not only is this model simpler and arguably easier to interpret, but it also constrains the parameter space to smoother and more manageable functions. In fact, it does include a model based on finite linear combinations of the marginals of $X$ as a particular case, while also generalizing the aforementioned $L^{2}$-models under some conditions. These RKHS-based models and their idiosyncrasies have been explored in Berrendero et al. (2019, 2024) in the linear setting, and in Berrendero et al. (2023) in the logistic case.
A major aim of this work is to motivate these models inside the functional framework, while also providing efficient techniques to apply them in practice. Our main contribution is the proposal of a Bayesian approach for inference in these RKHS models, in which a prior distribution is imposed on $\beta$ to use the posterior probabilities for prediction. Although setting a prior distribution on a functional space is generally a hard task, the specific parametric formulation we propose greatly facilitates this. Similar Bayesian schemes have recently been explored in Grollemund et al. (2019) and Abraham (2024), albeit not within a RKHS framework. Another set of techniques extensively studied in this context are variable selection methods, which aim to select the marginals $\{X(t_{j})\}$ of the process that better summarize it according to some optimality criterion (see Ferraty et al., 2010 or Berrendero et al., 2016 by way of illustration). As it happens, some RKHS-based variable selection methods have already been proposed (e.g. Bueno-Larraz and Klepsch, 2019), but in general they have their own dedicated algorithms and procedures. As will shortly become apparent, our Bayesian methodology allows us to easily isolate the marginal posterior distribution corresponding to a finite set of points $\{t_{j}\}$, providing a Bayesian variable selection process along with the other prediction methods that naturally arise from the posterior.
@@
The methodology proposed in this work relies heavily on the use of RKHS’s, so we outline the main characteristics of these spaces from a probabilistic point of view (for a more detailed account, see for example Berlinet and Thomas-Agnan, 2004). Let us denote by $K(t,s)=\mathbb{E}[X(t)X(s)]$ the covariance function of the centered process $X$, and in what follows suppose that it is continuous. To construct the corresponding RKHS $\mathcal{H}(K)$, we start by defining the functional space $\mathcal{H}_{0}(K)$ of all finite linear combinations of evaluations of $K$, that is,
$\mathcal{H}_{0}(K)=\left\{f\in L^{2}[0,1]:\ f(\cdot)=\sum_{j=1}^{p}a_{j}K(t_{j},\cdot),\ p\in\mathbb{N},\ a_{j}\in\mathbb{R},\ t_{j}\in[0,1]\right\}.$ (1.3)
@@
This space is endowed with the inner product $\left\langle f,g\right\rangle_{K}=\sum_{i,j}a_{i}b_{j}K(t_{i},s_{j})$, given that $f(\cdot)=\sum_{i}a_{i}K(t_{i},\cdot)$ and $g(\cdot)=\sum_{j}b_{j}K(s_{j},\cdot)$. Then, $\mathcal{H}(K)$ is defined to be the completion of $\mathcal{H}_{0}(K)$ under the norm induced by the scalar product $\left\langle\cdot,\cdot\right\rangle_{K}$. As it turns out, functions in this space satisfy the so-called reproducing property: $\left\langle K(t,\cdot),f\right\rangle_{K}=f(t)$ for all $f\in\mathcal{H}(K)$ and $t\in[0,1]$. An important consequence of this identity is that $\mathcal{H}(K)$ is a space of genuine functions and not of equivalence classes, since the values of the functions at specific points are in fact relevant, unlike in $L^{2}$-spaces.
@@
Now, a particularly useful approach in statistics is to regard $\mathcal{H}(K)$ as an isometric copy of a well-known space. Specifically, via Loève’s isometry (Loève, 1948) one can establish a congruence $\Psi_{X}$ between $\mathcal{H}(K)$ and the linear span of the process, $\mathcal{L}(X)$, in the space of all random variables with finite second moment, $L^{2}(\Omega)$ (see Lemma 1.1 in Lukić and Beder, 2001). This isometry is essentially the completion of the correspondence
$\sum_{j=1}^{p}a_{j}X(t_{j})\longleftrightarrow\sum_{j=1}^{p}a_{j}K(t_{j},\cdot),$
@@
and can be formally defined, in terms of its inverse, as $\Psi^{-1}_{X}(U)(t)=\mathbb{E}[UX(t)]$ for $U\in\mathcal{L}(X)$. Despite the close connection between the process $X$ and the space $\mathcal{H}(K)$, special care must be taken when dealing with concrete realizations, since in general the trajectories of $X$ do not belong to the corresponding RKHS with probability one (Lukić and Beder, 2001, Corollary 7.1). As a consequence, the expression $\left\langle x,f\right\rangle_{K}$ is ill-defined and lacks meaning when $x$ is a realization of $X$. However, following Parzen’s approach in his seminal work (Parzen, 1961, Theorem 4E), we can leverage Loève’s isometry and identify $\left\langle x,f\right\rangle_{K}$ with the image $\Psi_{x}(f):=\Psi_{X}(f)(\omega)$, for $x=X(\omega)$ and $f\in\mathcal{H}(K)$. This notation, viewed as a formal extension of the inner product, often proves to be useful and convenient.
@@
The rest of the paper is organized as follows. In Section 2 we explain the Bayesian methodology and the functional regression models we propose, including an overview of the reversible jump MCMC scheme. In Section 3 we derive a theoretical posterior consistency result. The empirical results of the experimentation are contained in Section 4, along with a short discussion of computational details. Lastly, the conclusions drawn from this work are presented in Section 5. Moreover, additional details and results are included in Appendices ABC and D.

### v3

@@ Abstract
We propose a novel Bayesian methodology for inference in functional linear and logistic regression models based on the theory of reproducing kernel Hilbert spaces (RKHS’s). We introduce general models that build upon the RKHS generated by the covariance function of the underlying stochastic process, and whose formulation includes as particular cases all finite-dimensional models based on linear combinations of marginals of the process, which can collectively be seen as a dense subspace made of simple approximations. By imposing a suitable prior distribution on this dense functional space we can perform data-driven inference via standard Bayes methodology, estimating the posterior distribution through reversible jump Markov chain Monte Carlo methods. In this context, our contribution is two-fold. First, we derive theoretical results that guarantee strong posterior consistency and contraction at an optimal rate under mild conditions. Second, we show that several prediction strategies stemming from our Bayesian procedure are competitive against other usual alternatives in both simulations and real data sets, including a Bayesian-motivated variable selection method.
@@ 1 Introduction
The problem of predicting a scalar response from a functional covariate is one that has gained traction over the last few decades, as more and more real-world data is being generated with an ever-increasing level of granularity in the measurements. While in principle the functional data could simply be regarded as a discretized vector in a very high dimension, there are often many advantages in taking into account its functional nature, ranging from modeling the correlation among points that are close in the domain, to extracting information that may be hidden in the derivatives of the function in question. As a consequence, numerous proposals have arisen on how to suitably deal with functional data, all of them encompassed under the term Functional Data Analysis (FDA), which essentially explores statistical techniques to process, model and make inference on data varying over a continuum. A partial survey on such methods is Cuevas (2014) or Goia and Vieu (2016), while a more detailed exposition of the theory and applications can be found for example in Hsing and Eubank (2015) and Horváth and Kokoszka (2012).
@@
FDA is undoubtedly an active area of research, which finds applications in a wide variety of fields, such as biomedicine, finance, meteorology or chemistry (Ullah and Finch, 2013). Accordingly, there are many recent contributions on how to tackle functional data problems, both from a theoretical and practical standpoint. Chief among them is the approach of reducing the problem to a finite-dimensional one, for example using a truncated basis expansion or spline interpolation methods (Müller and Stadtmüller, 2005; Aguilera and Aguilera-Morillo, 2013). At the same time, much effort has been put into building a sound theoretical basis for FDA, generalizing different frequentist concepts to the infinite-dimensional framework. Examples of this endeavor include the definition of centrality measures and depth-based notions for functional data (López-Pintado and Romo, 2009), functional ANOVA tests (Cuevas et al., 2004), a functional Mahalanobis distance (Galeano et al., 2015; Berrendero et al., 2020), or an extension of Fisher’s discriminant analysis for random functions (Shin, 2008), among many others. As the name suggests, FDA techniques are heavily inspired by functional analysis tools and methods: Hilbert spaces, linear operators, orthonormal bases, and so on. Incidentally, a notion that also intersects with the classical theory of machine learning and pattern recognition, and that has attained popularity in recent years, is that of reproducing kernel Hilbert spaces (RKHS’s) and their applications in functional data problems (Kupresanin et al., 2010; Yuan and Cai, 2010; Berrendero et al., 2018). On the other hand, Bayesian inference methods are ubiquitous in the realm of statistics, and though they also make use of random functions, the approach is slightly different from the FDA case. While there are recent works that offer a Bayesian treatment of functional data (e.g. Crainiceanu and Goldsmith, 2010; Shi and Choi, 2011), there is still no systematic approach to Bayesian methodologies within FDA. It is precisely at this relatively unexplored intersection between FDA and Bayesian methods that our work is aimed.
@@
In particular, our goal is to study functional regression problems, the infinite-dimensional equivalents of the typical statistical regression problems, from a Bayesian perspective. We follow the path started by Ferguson (1974) of setting a prior distribution on a functional space and using the corresponding posterior for inference. In our case, we use a particular RKHS as the ambient space, resulting in functional regression models that allow for a simple yet efficient Bayesian treatment of functional data. Moreover, we also study the basic theoretical question of posterior consistency in these RKHS models within the proposed Bayesian framework. Consistency and posterior concentration are a type of frequentist validation criteria that have arguably been an active point of research in the last few decades, particularly in infinite-dimensional settings (Amewou-Atisso et al., 2003; Choi and Ramamoorthi, 2008), and also in the functional regression case (Lian et al., 2016; Abraham and Grollemund, 2020). To put it simply, posterior consistency ensures that with enough data points, the Bayesian updating mechanism works as intended and the posterior distribution eventually concentrates around the true value of the parameters, supposing the model is well specified. We leverage the properties of RKHS’s and certain extensions of classical results by Doob (1949) and Schwartz (1965) to show that posterior consistency and contraction at an optimal rate hold in our models with minimal conditions, thus providing a coherent background to our Bayesian approach.
@@
Finally, this theoretical side is complemented by extensive experimentation that showcases the application of the functional RKHS models in various prediction tasks. Following recent trends in Bayesian computation techniques, the posterior distribution is approximated via Markov chain Monte Carlo (MCMC) methods, specifically the reversible jump variant (RJMCMC) proposed by Green (1995). This computational work highlights the predictive performance of the proposed functional regression models, especially when compared with other usual frequentist methods.
$L^{2}$-models, shortcomings and alternatives
@@
The most common scalar-on-function linear regression model is the classical $L^{2}$-model, widely popularized since the first edition (1997) of the pioneering monograph by Ramsay and Silverman (2005). It can be seen as a generalization of the usual finite-dimensional model, replacing the scalar product in $\mathbb{R}^{d}$ for that of the functional space $L^{2}[0,1]$ (henceforth denoted by $\left\langle\cdot,\cdot\right\rangle$):
$Y=\alpha+\left\langle X,\beta\right\rangle+\varepsilon=\alpha+\int_{0}^{1}X(t)\beta(t)\,dt+\varepsilon,$ (1.1)
where $\alpha\in\mathbb{R}$, $\varepsilon$ is a random error term independent from $X$ with $\mathbb{E}[\varepsilon]=0$, and the functional slope parameter $\beta=\beta(\cdot)$ is a member of the infinite-dimensional space $L^{2}[0,1]$. In this case, the inference on $\beta$ is hampered by the fact that $L^{2}[0,1]$ is an extremely broad space that contains many non-smooth or ill-behaved functions, so any estimation procedure involving optimization on it will typically be hard. In spite of this, model (1.1) is not flexible enough to include “simple” impact points models based on linear combinations of the marginals of $X$, such as $Y=\alpha+\beta_{1}X(t_{1})+\cdots+\beta_{p}X(t_{p})+\varepsilon$ for some constants $\beta_{j}\in\mathbb{R}$ and instants $t_{j}\in[0,1]$, which are especially appealing to practitioners confronted with functional data problems; see Berrendero et al. (2024) for additional details on this. Moreover, the non-invertibility of the covariance operator associated with $X$, which plays the role of the covariance matrix in the infinite case, invalidates the usual least squares theory (Cardot and Sarda, 2018). Thus, some regularization or dimensionality reduction technique is needed for parameter estimation; see Reiss et al. (2017) for a summary of several widespread methods.
@@
$\mathbb{P}(Y=1\mid X)=\frac{1}{1+\exp\{-\alpha-\left\langle X,\beta\right\rangle\}},$ (1.2)
where $\alpha\in\mathbb{R}$ and $\beta\in L^{2}[0,1]$. In the equivalent finite-dimensional problem the natural way of estimating the slope parameter is via its maximum likelihood estimator (MLE). However, apart from the issues outlined above for the linear model, functional settings pose an additional challenge in the logistic case: under fairly general conditions, the MLE does not exist with probability one (see Berrendero et al., 2023).
It turns out that in both scenarios a natural alternative to the $L^{2}$-model is the so-called reproducing kernel Hilbert space (RKHS) model, which instead assumes the unknown functional parameter $\beta$ to be a member of the RKHS associated with the covariance function of the process $X$. As we will show later on, not only is this model simpler and arguably easier to interpret, but it also constrains the parameter space to smoother and more manageable functions. In fact, it does include a model based on finite linear combinations of the marginals of $X$ as a particular case, while also generalizing the aforementioned $L^{2}$-models under some conditions. These RKHS-based models and their idiosyncrasies have been explored in Berrendero et al. (2019, 2024) in the linear case, and in Berrendero et al. (2023) in the logistic case.
A major aim of this work is to motivate these models within the functional framework, while also providing efficient techniques to apply them in practice. Our main contribution is the proposal of a Bayesian approach for inference in these RKHS models, in which a prior distribution is imposed on $\beta$ to use the posterior probabilities for prediction. Although placing a prior distribution on a functional space is generally a hard task, the specific parametric formulation we propose greatly facilitates this. Similar Bayesian schemes have recently been explored in Grollemund et al. (2019) and Abraham (2024), albeit not in a RKHS setting. Another set of techniques extensively studied in this context are variable selection methods, which aim to select the marginals $\{X(t_{j})\}$ of the process that better summarize it according to some optimality criterion (see Ferraty et al., 2010 or Berrendero et al., 2016 by way of illustration). As it happens, some RKHS-based variable selection methods have already been proposed (e.g. Bueno-Larraz and Klepsch, 2019), but in general they have their own dedicated algorithms and procedures. As will shortly become apparent, our Bayesian methodology allows us to easily isolate the marginal posterior distribution corresponding to a finite set of points $\{t_{j}\}$, providing a Bayesian variable selection process along with the other prediction methods that naturally arise from our approach.
@@
The methodology proposed in this work relies heavily on the use of RKHS’s, so we will briefly outline the main characteristics of these spaces from a probabilistic point of view (for a more detailed account, see Berlinet and Thomas-Agnan, 2004). Let us denote by $K(t,s)=\mathbb{E}[X(t)X(s)]$ the covariance function (the “kernel”) of the centered process $X$, and in what follows suppose that it is continuous. To construct the corresponding RKHS $\mathcal{H}(K)$, we start by defining the functional space $\mathcal{H}_{0}(K)$ of all finite linear combinations of evaluations of $K$, that is,
$\mathcal{H}_{0}(K)=\left\{f\in L^{2}[0,1]:\ f(\cdot)=\sum_{j=1}^{p}\beta_{j}K(t_{j},\cdot),\ p\in\mathbb{N},\ \beta_{j}\in\mathbb{R},\ t_{j}\in[0,1]\right\}.$ (1.3)
@@
This space is endowed with the inner product $\left\langle f,g\right\rangle_{K}=\sum_{j,k}\beta_{j}\gamma_{k}K(t_{j},s_{k})$, given that $f(\cdot)=\sum_{j}\beta_{j}K(t_{j},\cdot)$ and $g(\cdot)=\sum_{k}\gamma_{k}K(s_{k},\cdot)$. Then, $\mathcal{H}(K)$ is defined to be the completion of $\mathcal{H}_{0}(K)$ under the norm induced by the scalar product $\left\langle\cdot,\cdot\right\rangle_{K}$. As it turns out, functions in this space satisfy the so-called reproducing property: $\left\langle K(t,\cdot),f\right\rangle_{K}=f(t)$ for all $f\in\mathcal{H}(K)$ and $t\in[0,1]$. An important consequence of this identity is that $\mathcal{H}(K)$ is a space of genuine functions and not of equivalence classes, since the values of the functions at specific points are in fact relevant, unlike in $L^{2}$-spaces.
@@
Now, a particularly useful approach in statistics is to regard $\mathcal{H}(K)$ as an isometric copy of a well-known space related to $X$. Specifically, via Loève’s isometry (Loève, 1948) one can establish a congruence $\Psi_{X}$ between $\mathcal{H}(K)$ and the linear span of the process, $\mathcal{L}(X)$, in the space of all random variables with finite second moment, $L^{2}(\Omega)$ (see Lemma 1.1 in Lukić and Beder, 2001). This isometry is essentially the completion of the correspondence
$\sum_{j=1}^{p}\beta_{j}X(t_{j})\longleftrightarrow\sum_{j=1}^{p}\beta_{j}K(t_{j},\cdot),$
@@
and can be formally defined, in terms of its inverse, as $\Psi^{-1}_{X}(U)(t)=\mathbb{E}[UX(t)]$ for $U\in\mathcal{L}(X)$. Despite the close connection between the process $X$ and the space $\mathcal{H}(K)$, special care must be taken when dealing with concrete realizations, since in general the trajectories of $X$ do not belong to the corresponding RKHS with probability one (Lukić and Beder, 2001, Corollary 7.1). As a consequence, the expression $\left\langle x,f\right\rangle_{K}$ is ill-defined and lacks meaning when $x$ is a realization of $X$. However, following Parzen’s approach in his seminal work (Parzen, 1961, Theorem 4E), we can leverage Loève’s isometry and identify $\left\langle X,f\right\rangle_{K}$ with the random variable in $\mathcal{L}(X)$ associated with $f\in\mathcal{H}(K)$, so that $\left\langle x,f\right\rangle_{K}$ with $x=X(\omega)$ is defined as the image $\Psi_{x}(f):=\Psi_{X}(f)(\omega)$. This notation, viewed as a formal extension of the inner product, often proves to be useful and convenient.
@@
The rest of the paper is organized as follows. In Section 2 we explain the Bayesian methodology and the functional regression models we propose, including an overview of the reversible jump MCMC scheme. In Section 3 we derive theoretical posterior consistency and posterior concentration results. The empirical findings of the experimentation are contained in Section 4, along with a short discussion of computational details. Lastly, the conclusions drawn from this work are presented in Section 5. Additional details, proofs and results are included in Appendices A, B, C and D.